OpenStudy (anonymous):

Solve using elimination-by-addition method: 9x -6y = 48 10x + 5y = 30

OpenStudy (anonymous):

Oh! I think I solved it! :D

OpenStudy (ash2326):

we \[9x-6y=48\] and \[10x+5y=30\] let's eliminate y from the two, let's multiply the first one by 10 and second equation by 12 we have \[10(9x-6y)=10*48\] and \[12(10x+5y)=12*30\] so we get \[90x-60y=480\] and \[120x+60y=360\] let's add the two equations, this will eliminate y \[90x+120x=480+360\] now combining like terms we get \[210x=840\] dividing both sides by 210 we get \[x=4\] now to find y let's substitute x in original first equation (We can use any equation to determine y) \[9x-6y=48\] x=4 \[9*4-6y=48\] \[36-6y=48\] subtracting both sides by 6 we get \[-6y=12\] dividing both sides by -2 , we get \[y=-2\] so x=4 and y =-2

OpenStudy (ash2326):

Good Mandi, I didn't see your post, and continued solving. :$

OpenStudy (anonymous):

THANK YOU!